OpenStudy (anonymous):

Simplify showing your work. Remember to use a ^ symbol to indicate exponents. (2bc^3)^2 (-3b) =

OpenStudy (anonymous):

The first step is to distribute ^2 to everything inside the parenthesis

OpenStudy (fibonaccichick666):

...whoelse you should rephrase that

OpenStudy (anonymous):

You want to raise everything to the power of two that is inside the parenthesis by distribution

OpenStudy (fibonaccichick666):

ok I am content with that

OpenStudy (unklerhaukus):

\[(2bc^3)^2 (-3b) =(2bccc)(2bccc) (-3b) =\quad \dots\]

OpenStudy (anonymous):

not sure

OpenStudy (unklerhaukus):

\[=(2\times2\times-3)bbbcccccc=\quad\dots\]

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

2^2 =? b^2 =? (c^3)^2 =?